### v1

@@ 5.2 Reliability of the derived burst energy
The burst energy estimate was obtained using two approaches: we compared the HAWC+ post-burst flux density ratios and the Ks ratio curve individually (for all three settings). The results agree pretty well, which strengthens our approach. Our final estimate of the burst energy is based on the fit to the ratio curve, while the confidence intervals are based on the HAWC+ fits.
In the following, we will discuss how well the flux increase is reproduced by our models. We neglect all the changes that may occur during the burst. But this will not be the case in reality. The most obvious change is probably the shift of the sublimation radius due to the burst. To investigate the effect, we performed a simulation that includes dust sublimation. We set the innermost radius at (20 au) and chose a time step of 0.7 days. Dust with a temperature greater than 1 600 K was removed at each time step. The radiation is produced by the hottest dust. Therefore, we cut the grid at the outer edge of the disk. With this choice, the resolution in the innermost regions is higher, while the entire flux is captured. The result is plotted in Fig. 22. The maximum ratio is about a factor of 1.5 smaller for the more realistic sublimation model (blue) compared to the mean model (black). The burst energy (and shape) are the same in both cases. This result might be surprising, as the emitting surface area should increase as the inner rim moves outward. Therefore, the ratio should be higher when dust sublimation is included. However, the inner radius of the mean model was set at to avoid unrealistic high temperatures. Thus, the area where the -band photons are emitted is larger. This shows that the most probable value (for ) is slightly higher than our estimate, which does neglect dust sublimation. One could argue that a small fraction of the burst energy is needed to sublimate the disk, which might compensate for the above effect. However, this fraction is probably negligible. Typical values for the sublimation enthalpy range from 1 (hydrogen) to almost (graphite). With a molar mass of () and a gas-to-dust ratio of 100, it would take less than to sublimate (on the order of a total disk mass or more). This is less than of the burst energy.
@@ 5.3 What was swallowed by G323?
The properties of the HCHii region can be used to assess the evolutionary status of G323. The free-free emission of most hyper/ultra-compact Hii regions associated with MYSOs points to an excess of Lyman photons over pure photospheric emission, which is attributed to accretion (Cesaroni et al., 2015, 2016). For G323, the Lyman continuum photon flux can be derived using Eq. 10 in Martín-Hernández et al. (2005) by plugging in the 3 mm flux density of Jy measured by us, the electron temperature of K obtained by Zhang et al. (2023), and the distance given above. The resulting value of is lower than the prediction of from the relation for ZAMS stars (Cesaroni et al., 2016) for the luminosity range of our models. The discrepancy becomes even greater when a possible contribution of an accretion shock is taken into account. This suggests that G323 is still in the pre-ZAMS state at a lower effective temperature, making the bloating scenario seem feasible.
Together, we conclude that the swallowed object was very likely much heavier than our estimate given in Sect. 4.4.3.
@@ 5.5 The inner region and the protostellar disk
Note that during the burst, the sublimation radius is expected to shift outward by a few 10 au. The -band observation was done nearly a decade after the peak of the burst. This is too short for the material to recondense. Thus, the inner rim at the date of the -band observations may be further away than . Indeed, for low-mass YSO bursts, dust sublimation fronts are an indirect tracer of past outbursts (see, e.g., Jørgensen et al., 2020). However, for massive protostars, the matter will be replenished quickly, given the high accretion rates. With an accretion rate of , it follows that in a decade can be transported to the inner region, enough to refill the inner disk. If the sublimation radius were still further out, higher fluxes would result from the increased surface area. However, the -band post- and pre-burst fluxes are the same within the errors. Together, this indicates that there are no significant changes present in the dust distribution in 2023 and that the -band emission originates at the inner rim of the disk, which seems to be as close as the pre-burst dust sublimation radius.
@@ 5.6 Prospects for TDRT modeling of MYSO bursts
The present paper describes the first application of TDRT based on the TORUS code in a real-world case. In that context, we want to outline its potential for the future.
The time-dependent approach we use is only the first step in the direction of a fully self-consistent treatment of outbursts. Improved and more systematic modeling is planned for existing FIR bursts (including G323). The next step is to use larger sets of time-dependent models, which all agree with the pre-burst. With this in mind, the previous parameter ranges (for both burst and MYSO geometry) shall be refined in the near future.
Due to the SOFIA shutdown, no further FIR observations will be possible in the mid-term. However, future MYSO bursts can be studied with ground-based facilities in the (sub)mm and in the NIR/MIR provided the burst hosts are no longer deeply embedded. JWST is capable of targeting younger and more enshrouded objects.
With TORUS the dust continuum can be modeled in those spectral ranges as well, while line emission is currently not included in the time-dependent version. The temperature grids are delivered at each time step. They might be used as input for chemical burst models, similar to the approach used in Rab et al. (2017); Guadarrama et al. (2024). TORUS does not include chemical networks, which are needed to directly obtain chemical burst models.
Another application of the temperature output is to constrain the location of possible methanol maser sites. This was done in a simplified manner by Stecklum et al. (2021) for static model grids.
For all our models, we used very simplified assumptions, namely no protostellar bloating, axis-symmetry, same dust in all regions and no changes during the burst besides heating and cooling (i.e., no changes in the dust distribution and chemistry due to sublimation, etc.). Some of these limitations might be explored in the future. A simple form of dust sublimation is included in which a fraction of the dust exceeding the sublimation temperature is removed each time step. Currently, resublimation is immediate once the temperature drops below the sublimation temperature. This might be improved in the future.
Eventually, we emphasize that TDRT can be applied to other classes of variable objects and transient phenomena where dust plays a major role.
@@ 6 Conclusions
G323 featured a powerful accretion burst that peaked in 2013. It extends the small sample of known MYSO bursts. The objective of this work is to draw conclusions about the nature of the burst from the derived limits for the released energy and the accreted mass. G323’s burst is a multiwavelength phenomenon that was accompanied by flares of different maser species. It was observed in the NIR, MIR, FIR, and radio at different timescales.
There is a clear correlation between the Class II methanol maser flare and the IR radiation. A light echo is present in , and as well as in the images. Interestingly, its expansion appears to be faster at longer wavelengths, possibly because at these wavelengths the scattering optical depth is reduced.
@@
Our SOFIA/HAWC+ observations, performed in 2022, two years after the end of the burst, constrained the strength of the thermal afterglow and were crucial to derive limits on the burst energy. These measurements and the light curve from the VVV(X) survey, indicate that the burst is probably the most energetic of all known MYSO bursts. The energy released is equal to . Its duration, i.e, , is in line with the accretion of a compact object. This is in accordance with the observed bluening during the burst, which is not expected if the material flow arises in a more diffuse stream from an active disk. For a protostar close to the ZAMS, the minimum mass needed to release this amount of energy is times the Jupiter mass, which is in the range of a big Jupiter, a brown dwarf, or a disk fragment.
The pre-burst luminosity of G323 is consistent with that of a bloated protostar, which might feature pulsations with the same period as the maser. If the bloating scenario holds, the accreted mass may be as high as half a solar mass.
The post-burst emission in the -band can be attributed to the inner disk region. It indicates a minimum extent of the emitting region of 14 mas or 57 au, which is consistent with a disk inner rim situated at the dust sublimation radius during quiescence.
@@
For the first time, we use time-dependent radiative transfer (TDRT) to study a real science case. With TDRT, we can model the timescales self-consistently. The thermal afterglow is much longer than the grid crossing time. G323 is most likely seen face-on. Therefore, the NIR timescales are equal to those of the accretion rate variation. The light curve reflects the protostellar luminosity variation throughout the entire burst duration. At longer wavelengths, the timescales are much longer. We find that the visibility of the (MIR/FIR) afterglow can vary by up to years, depending on the dust distribution and the burst characteristics. The example of G323 shows that TDRT simulations are a powerful tool for investigating transient phenomena of dusty objects. For episodic accretion of MYSO, it opens up the possibility to obtain reliable burst parameters, which are needed to understand the unsteady protostellar growth, especially in the high-mass regime. The shutdown of SOFIA is a serious drawback, since it was the only facility that could observe in the FIR which is crucial to derive the burst energy. Future studies must focus on MIR and (sub)mm.
The rare face-on view of G323 and its NIR record of the burst make this source a key object for studying the consequences of episodic accretion and a benchmark for the corresponding simulations.
@@ Acknowledgements.
We thank the anonymous referee for accurate review and highly appreciate comments and suggestions which helped to improve the quality of this paper. The authors thank Dr. Gordon MacLeod for support concerning the maser observations and Dr. Riccardo Cesaroni for input concerning the free-free emission luminosity. VW was supported by the German Aerospace Center (DLR) under grant number 50OR1718. A.C.G. acknowledges from PRIN-MUR 2022 20228JPA3A ”The path to star and planet formation in the JWST era (PATH)” funded by NextGeneration EU and by INAF-GoG 2022 ”NIR-dark Accretion Outbursts in Massive Young Stellar Objects (NAOMY)” and Large Grant INAF 2022 ”YSOs, Outflows, Disks, and Accretion: towards a global framework for the evolution of planet-forming systems (YODA)”. This publication uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer ((NEO)WISE), which is a joint project of the Jet Propulsion Laboratory/California Institute of Technology and the University of Arizona. (NEO)WISE is funded by the National Aeronautics and Space Administration. Based on data products from observations made with ESO Telescopes at the La Silla or Paranal Observatories under ESO programmes ID083.C-0582, ID 179.B-2002 and 198.B-2004. Based on data obtained from the ESO Science Archive Facility under request numbers 557557 and 573746. The ALMA data are from the programme ALMAGAL (2019.1.00195.L) and were retrieved from the ALMA Science Archive (e.g., Stoehr et al., 2014). This paper used information from the Red MSX Source survey database at http://rms.leeds.ac.uk/cgi-bin/public/RMS_DATABASE.cgi, which was constructed with the support of the Science and Technology Facilities Council of the UK. This work used data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. The national facility capability for SkyMapper has been funded through ARC LIEF grant LE130100104 from the Australian Research Council, awarded to the University of Sydney, the Australian National University, Swinburne University of Technology, the University of Queensland, the University of Western Australia, the University of Melbourne, Curtin University of Technology, Monash University and the Australian Astronomical Observatory. SkyMapper is owned and operated by The Australian National University’s Research School of Astronomy and Astrophysics. Survey data were processed and provided by the SkyMapper Team at ANU. This research has made use of NASA’s Astrophysics Data System. This research has used adstex (https://github.com/yymao/adstex).References

### v2

@@ 5.2 Reliability of the derived burst energy
The burst energy estimate was obtained using two approaches: we compared the HAWC+ post-burst flux density ratios and the Ks ratio curve individually (for all three settings). The results agree pretty well, which strengthens our approach. Our final estimate of the burst energy is based on the fit to the ratio curve, while the confidence intervals are based on the HAWC+ fits. In the following, we discuss how well the flux increase is reproduced by our models. We neglect all the changes that may occur during the burst. But this will not be the case in reality. The most obvious change is probably the shift of the sublimation radius due to the burst. To investigate the effect, we performed a simulation that includes dust sublimation. We set the innermost radius at (20 au) and chose a time step of 0.7 days. Dust with a temperature greater than 1 600 K was removed at each time step. The radiation is produced by the hottest dust. Therefore, we cut the grid at the outer edge of the disk. With this choice, the resolution in the innermost regions is higher, while the entire flux is captured. The result is plotted in Fig. 22. The maximum ratio is about a factor of 1.5 smaller for the more realistic sublimation model (blue) compared to the mean model (black). The burst energy (and shape) are the same in both cases. This result might be surprising, as the emitting surface area should increase as the inner rim moves outward. Therefore, the ratio should be higher when dust sublimation is included. However, the inner radius of the mean model was set at to avoid unrealistic high temperatures. Thus, the area where the -band photons are emitted is larger. This shows that the most probable value (for ) is slightly higher than our estimate, which does neglect dust sublimation. One could argue that a small fraction of the burst energy is needed to sublimate the disk, which might compensate for the above effect. However, this fraction is probably negligible. Typical values for the sublimation enthalpy range from 1 (hydrogen) to almost (graphite). With a molar mass of () and a gas-to-dust ratio of 100, it would take less than to sublimate (on the order of a total disk mass or more). This is less than of the burst energy.
@@ 5.3 The accreted object
The properties of the HCHii region can be used to assess the evolutionary status of G323. The free-free emission of most hyper/ultra-compact Hii regions associated with MYSOs points to an excess of Lyman photons over pure photospheric emission, which is attributed to accretion (Cesaroni et al., 2015, 2016). For G323, the Lyman continuum photon flux can be derived using Eq. 10 in Martín-Hernández et al. (2005) by plugging in the 3 mm flux density of Jy measured by us, the electron temperature of K obtained by Zhang et al. (2023), and the distance given above. The resulting value of is lower than the prediction of from the relation for ZAMS stars (Cesaroni et al., 2016) for the luminosity range of our models. The discrepancy becomes even greater when a possible contribution of an accretion shock is taken into account. This suggests that G323 is still in the pre-ZAMS state at a lower effective temperature, making the bloating scenario seem feasible. Together, we conclude that the swallowed object was very likely much heavier than our estimate given in Sect. 4.4.3.
@@ 5.5 The inner region and the protostellar disk
We note that during the burst, the sublimation radius is expected to shift outward by a few 10 au. The -band observation was done nearly a decade after the peak of the burst. This is too short for the material to recondense. Thus, the inner rim at the date of the -band observations may be further away than . Indeed, for low-mass YSO bursts, dust sublimation fronts are an indirect tracer of past outbursts (see, e.g., Jørgensen et al., 2020). However, for massive protostars, the matter will be replenished quickly, given the high accretion rates. With an accretion rate of , it follows that in a decade can be transported to the inner region, enough to refill the inner disk. If the sublimation radius were still further out, higher fluxes would result from the increased surface area. However, the -band post- and pre-burst fluxes are the same within the errors. Together, this indicates that there are no significant changes present in the dust distribution in 2023 and that the -band emission originates at the inner rim of the disk, which seems to be as close as the pre-burst dust sublimation radius.
@@ 5.6 Prospects for TDRT modeling of MYSO bursts
The present paper describes the first application of TDRT based on the TORUS code in a real-world case. In that context, we want to outline its potential for the future. The time-dependent approach we use is only the first step in the direction of a fully self-consistent treatment of outbursts. Improved and more systematic modeling is planned for existing FIR bursts (including G323). The next step is to use larger sets of time-dependent models, which all agree with the pre-burst. With this in mind, the previous parameter ranges (for both burst and MYSO geometry) shall be refined in the near future. Due to the SOFIA shutdown, no further FIR observations will be possible in the mid-term. However, future MYSO bursts can be studied with ground-based facilities in the (sub)mm and in the NIR/MIR provided the burst hosts are no longer deeply embedded. JWST is capable of targeting younger and more enshrouded objects.
With TORUS the dust continuum can be modeled in those spectral ranges as well, while line emission is currently not included in the time-dependent version. The temperature grids are delivered at each time step. They might be used as input for chemical burst models, similar to the approach used in Rab et al. (2017); Guadarrama et al. (2024). TORUS does not include chemical networks, which are needed to directly obtain chemical burst models. Another application of the temperature output is to constrain the location of possible methanol maser sites. This was done in a simplified manner by Stecklum et al. (2021) for static model grids.
For all our models, we used very simplified assumptions, namely no protostellar bloating, axis-symmetry, same dust in all regions and no changes during the burst besides heating and cooling (i.e., no changes in the dust distribution and chemistry due to sublimation, etc.). Some of these limitations might be explored in the future. A simple form of dust sublimation is included in which a fraction of the dust exceeding the sublimation temperature is removed each time step. Currently, resublimation is immediate once the temperature drops below the sublimation temperature. This might be improved in the future. Eventually, we emphasize that TDRT can be applied to other classes of variable objects and transient phenomena where dust plays a major role.
@@ 6 Conclusions
G323 featured a powerful accretion burst that peaked in 2013. It extends the small sample of known MYSO bursts. The objective of this work is to draw conclusions about the nature of the burst from the derived limits for the released energy and the accreted mass. G323’s burst is a multiwavelength phenomenon that was accompanied by flares of different maser species. It was observed in the NIR, MIR, FIR, and radio at different timescales. There is a clear correlation between the Class II methanol maser flare and the IR radiation. A light echo is present in , and as well as in the images. Interestingly, its expansion appears to be faster at longer wavelengths, possibly because at these wavelengths the scattering optical depth is reduced.
@@
Our SOFIA/HAWC+ observations, performed in 2022, two years after the end of the burst, constrained the strength of the thermal afterglow and were crucial to derive limits on the burst energy. These measurements and the light curve from the VVV(X) survey, indicate that the burst is probably the most energetic of all known MYSO bursts. The energy released is equal to . Its duration, i.e, , is in line with the accretion of a compact object. This is in accordance with the observed bluening during the burst, which is not expected if the material flow arises in a more diffuse stream from an active disk. For a protostar close to the ZAMS, the minimum mass needed to release this amount of energy is times the Jupiter mass, which is in the range of a big Jupiter, a brown dwarf, or a disk fragment. The pre-burst luminosity of G323 is consistent with that of a bloated protostar, which might feature pulsations with the same period as the maser. If the bloating scenario holds, the accreted mass may be as high as half a solar mass. The post-burst emission in the -band can be attributed to the inner disk region. It indicates a minimum extent of the emitting region of 14 mas or 57 au, which is consistent with a disk inner rim situated at the dust sublimation radius during quiescence.
@@
For the first time, we used time-dependent radiative transfer (TDRT) to study a real science case. With TDRT, we can model the timescales self-consistently. The thermal afterglow is much longer than the grid crossing time. G323 is most likely seen face-on. Therefore, the NIR timescales are equal to those of the accretion rate variation. The light curve reflects the protostellar luminosity variation throughout the entire burst duration. At longer wavelengths, the timescales are much longer. We find that the visibility of the (MIR/FIR) afterglow can vary by up to years, depending on the dust distribution and the burst characteristics. The example of G323 shows that TDRT simulations are a powerful tool for investigating transient phenomena of dusty objects. For episodic accretion of MYSO, it opens up the possibility to obtain reliable burst parameters, which are needed to understand the unsteady protostellar growth, especially in the high-mass regime. The shutdown of SOFIA is a serious drawback, since it was the only facility that could observe in the FIR which is crucial to derive the burst energy. Future studies must focus on MIR and (sub)mm. The rare face-on view of G323 and its NIR record of the burst make this source a key object for studying the consequences of episodic accretion and a benchmark for the corresponding simulations.
@@ Acknowledgements.
We thank the anonymous referee for accurate review and highly appreciate comments and suggestions which helped to improve the quality of this paper. The authors thank Dr. Gordon MacLeod for support concerning the maser observations and Dr. Riccardo Cesaroni for input concerning the free-free emission luminosity. VW was supported by the German Aerospace Center (DLR) under grant number 50OR1718. A.C.G. acknowledges from PRIN-MUR 2022 20228JPA3A ”The path to star and planet formation in the JWST era (PATH)” funded by NextGeneration EU and by INAF-GoG 2022 ”NIR-dark Accretion Outbursts in Massive Young Stellar Objects (NAOMY)” and Large Grant INAF 2022 ”YSOs, Outflows, Disks, and Accretion: toward a global framework for the evolution of planet-forming systems (YODA)”. This publication uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer ((NEO)WISE), which is a joint project of the Jet Propulsion Laboratory/California Institute of Technology and the University of Arizona. (NEO)WISE is funded by the National Aeronautics and Space Administration. Based on data products from observations made with ESO Telescopes at the La Silla or Paranal Observatories under ESO programmes ID083.C-0582, ID 179.B-2002 and 198.B-2004. Based on data obtained from the ESO Science Archive Facility under request numbers 557557 and 573746. The ALMA data are from the programme ALMAGAL (2019.1.00195.L) and were retrieved from the ALMA Science Archive (e.g., Stoehr et al., 2014). This paper used information from the Red MSX Source survey database at http://rms.leeds.ac.uk/cgi-bin/public/RMS_DATABASE.cgi, which was constructed with the support of the Science and Technology Facilities Council of the UK. This work used data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. The national facility capability for SkyMapper has been funded through ARC LIEF grant LE130100104 from the Australian Research Council, awarded to the University of Sydney, the Australian National University, Swinburne University of Technology, the University of Queensland, the University of Western Australia, the University of Melbourne, Curtin University of Technology, Monash University and the Australian Astronomical Observatory. SkyMapper is owned and operated by The Australian National University’s Research School of Astronomy and Astrophysics. Survey data were processed and provided by the SkyMapper Team at ANU. This research has made use of NASA’s Astrophysics Data System. This research has used adstex (https://github.com/yymao/adstex).References